Optima Company is a high-technology organisation that produces a mass-storage system. The design of Optimas system is unique and represents a breakthrough in the industry. The units Optima produces combine positive features of both compact and hard disks. The company is completing its fifth year of operations and is preparing to build its master budget for the coming year (2017). The budget will detail each quarters activity and the activity for the year in total. The master budget will be based on the following information: a Fourth-quarter sales for 2016 are 65 000 units. b Unit sales by quarter [for 2017] are projected as follows: 0 First quarter Second quarter Third quarter Fourth quarter 75000 80000 85000 95000 The selling price is $500 per unit. All sale5 are credit sales. Optima collects 85% of all sales within the quarter in which they are realised; the other 15% is collected in the following quarter. There are no bad debts There is no beginning inventory of finished goods, Optima is planning the following ending finished goods inventories for each quarter c First quarter Second quarter Third quarter Fourth quarter 13000 units 15000 units 20000 units 10000 units d Each mass-storage unit uses five hours of direct labour and three units of direct materials. Workers are paid $25 per hour, and one unit of direct materials costs $80. There are 65 700 units of direct materials in beginning inventory as at 1 January 2017. At the end of each quarter, Optima plans to have 30% of the direct materials needed for next quarters unit sales. Optima will end the year with the same amount of direct materials found in this years beginning e f Optima buys direct materials on account. Half of the purchases are paid for in the quarter of acquisition and the remaining half are paid for in the following quarter. Wages and salaries are paid on the 15th and 30th of each month. Fixed overhead totals $1 million each quarter. Of this total, $350000 represents depreciation. All other fixed expenses are paid for in cash in the quarter incurred. The fixed overhead rate is computed by dividing the years total fixed overhead by the years budgeted production in units. g Variable overhead is budgeted at $6 per direct labour hour. All variable overhead expenses are paid for in the quarter incurred Fixed selling and administrative expenses total $250 000 per quarter, including $50 000 depreciation. Variable selling and administrative expenses are budgeted at $10 per unit sold. All selling and administrative expenses are paid for in the quarter incurred The balance sheet as at 31 December 2017 is as follows: h i j k
media%2Fba1%2Fba16bac7-c1a6-4052-857b-10
Answer #1

GIVEN DATA:

Fourth-quarter sales for 2016 = 65 000 units.

First quarter = 75 000
Second quarter = 80 000
Third quarter = 85 000
Fourth quarter = 95 000

Selling price = $500 per unit

Optima collects = 85% of all sales

other  collected = 15%

Ending finished goods inventories for each quarter:

First quarter = 13 000 units
Second quarter = 15 000 units
Third quarter = 20 000 units
Fourth quarter = 10 000 units

Direct materials costs = $80.

Direct materials in beginning inventory = 65700 units

And some additional information & balance sheet details as shown in the above information ect...

REQUIRTED:

1. Sales budget

2. Production budget

3. Direct materials purchases budget

4. Direct labor budget

5. Overhead budget

6. Selling and administrative expenses budget

7. Ending finished goods inventory budget

8. Cost of goods sold budget (Assume that there is no change in work-in-process

inventories.)

9. Cash budget

10. Pro forma income statement (using absorption costing)

11. Pro forma balance sheet

Expert Answer

SOLUTION:

1. OPTIMA COMPANY SALES BUDGET :

PARTICULARS QUARTER 1 QUARTER 2 QUARTER 3 QUARTER 4 TOTAL Units sold 75000 95000 335000 Sale price /unit Total sales(S) 37500000$ 40000000$ 42500000 47500000$ 167500000$

WORKING NOTE:

unit sold * sale price = Total sales

Q1 : 75000 *500 = 37500000$

Q2 : 80000 * 500 = 40000000$

Q3 :85000 * 500 = 42500000$

Q4: 95000 * 500 = 47500000$

There fore total sales budget is 167500000$

2. OPTIMA COMPANY PRODUCTION BUDGET:

WORKING NOTE:

PRODUCTION UNITS : sales units + closing inventory = total requiredment - opening stock.

Q1 : 75000 +13000 = 88000 - NIL  

= 88000

remaining Q2, Q3, Q4 Also calculate like that only

There fore total Production budget is 345000$

3. OPTIMA COMPANY DIRECT MATERIALS PURCHASE BUDGET:

4.OPTIMA COMPANY DIRECT LOBOR BUDGET:

Note : I have subject knowledge for above questions due to insufficent time ican't answer remaining questions please, post as another question